# soft open fences

### v1

@@ Abstract
We trained 13,440 large language models and found that entropy minimization requires only a single unlabeled data to achieve performance improvements comparable to or even greater than those obtained using thousands of data and carefully designed rewards in rule-based reinforcement learning. This striking result may prompt a rethinking of post-training paradigms for large language models. Our code is avaliable at https://github.com/zitian-gao/one-shot-em.
@@ 1 Introduction
The post-training phase of large language models (LLMs) has advanced rapidly [18, 17, 7, 16, 2, 20], with models like DeepSeek-R1 [3], Kimi-K1.5 [15], and OpenAI-o1 [10] demonstrating remarkable reasoning abilities. However, preparing for reinforcement learning is never an easy task. It often requires a large amount of high-quality ground truth labeled data, along with the careful design of strict rule-based reward functions to maximize advantage signals and prevent reward hacking. In contrast, entropy minimization is entirely unsupervised. We demonstrate, by training 13,440 large language models using entropy minimization, that with just one piece of unlabeled data, models trained in this way can achieve performance comparable to or even better than traditional reinforcement learning. Moreover, they typically converge within just 10 training steps, which is significantly faster than the thousands of steps often required for reinforcement learning. Entropy minimization is based on two direct and simple assumptions: (1) The generation and sampling process of large language models is inherently stochastic, (2) Correct answers generally have lower entropy than incorrect ones, which is similar to the optimization assumption used in pretraining. Based on our detailed study, we find that both Entropy Minimization (EM) and Reinforcement Learning (RL) share a common goal: unlocking the latent potential acquired during pretraining. Both methods operate on the same principle—introducing no new knowledge or capabilities, but instead maximizing the utility of the pretrained model, a process we refer to as token reranking. Our comprehensive investigation into the potential of entropy minimization reveals that it has the capacity to rival reinforcement learning in the post-training phase aimed at enhancing reasoning ability.
@@
  • We conduct an in-depth analysis of the effectiveness of One-shot EM, making it a highly reasonable approach. We find that it shares many core properties with RL, yet drives model behavior in the opposite direction when viewed through the lens of logit shift.
@@ 2.2 Data Selection
Entropy Minimisation (EM) relies on the premise that the model’s predictive uncertainty can serve as a meaningful training signal. However, not all input prompts are equally informative in this regard. As noted in prior work [12, 16], certain prompts elicit deterministic behavior from the model (e.g., always correct or always incorrect), yielding limited gradient information under entropy-based objectives.
@@
This variance quantifies the inconsistency of the model’s predictions for a given input. A low variance indicates either high confidence in correctness (near-perfect success) or high confidence in failure (uniformly wrong), both of which are suboptimal for entropy minimisation, as they lead to low-entropy posteriors that cannot be further improved.
@@
Intuitively, data with high pass@k variance suggests that the model’s response distribution is straddling the decision boundary—sometimes correct, sometimes not—indicating a broad or multimodal predictive distribution. These are precisely the regions where entropy minimisation is most impactful: it encourages the model to concentrate its probability mass on a consistent and (ideally) correct reasoning trajectory.
@@ 3.1 Experimental Setting
We implemented the overall training process of EM based on Acclerate [5]. We selected 1 piece of data from the dataset as prompt. Since it is an unsupervised method, we do not need any data labels. We directly train the model for only 10 steps with a constant learning rate of $2\times 10^{-5}$, a temperature parameter of 0.5, and a batch size of 64.  [^1] The reason why only 10 steps are sufficient will be explained in detail in [1].
@@ 3.2 Main Result
During training, the model repeatedly reinforces its own confidence, causing the logits distribution to gradually shift to the right—a phenomenon that will also be discussed in detail in [1].Compared to the initial average length of 500 tokens, after 1k steps of RL, the output has almost linearly and steadily increased to 2000 tokens, a significant increase of 4 times. As the response length increases, the model begins to exhibit more complex behaviors, such as reflection and exploration of alternative solutions. These behaviors emerge naturally, with no related data in our training set, and enhance the model’s ability to handle more complex tasks. These phenomena align closely with the results of R1 [3]. We will further discuss this in Section 3.3.
Temperature Batch Size Learning Rate Max Response Len
0.5 64 2e-5 2048
Table 2: Training Hyper-parameters
@@ 3.3 Logits Shift
We sample 20 prompts from the NuminaMath dataset and generate responses using four different models (Qwen2.5-Math-7B, Qwen2.5-Math-7B-EM, Qwen2.5-Math-7B-RL, Qwen2.5-Math-7B-EM-RL). Each model generates 20 responses, resulting in a total of $4\times 20=80$ outputs. For each output, we extract the logits, defined as the unnormalized scores $z_{i}\in\mathbb{R}$ produced by the model before applying the softmax function.
@@
In contrast, models trained with RL show a pronounced reduction in logits skewness, indicating a leftward shift in the distribution. We hypothesize that this is due to the influence of ground-truth signals during RL training. Specifically, RL appears to perform re-ranking that suppresses tokens with high predicted probability but low alignment with ground truth, thereby reducing their rank and shifting the overall distribution leftward.
@@
Figure 2: The left y-axis represents the EM training loss, while the right y-axis shows the average score across six reasoning benchmarks. It can be observed that the training loss converges rapidly, whereas the average score peaks around step 10 and then begins to decline. The results in the figure are obtained by repeating the experiments with the same training parameters using 16 different seeds to reduce randomness.
As shown in Figure 2, the loss drops to a relatively low level around step 10, and the model’s performance on mathematical reasoning reaches its peak. However, unexpectedly, as the EM training loss continues to decrease beyond step 10, the mathematical reasoning performance begins to decline. In conjunction with the findings presented in Section [2], we argue that EM functions primarily as a tool for shaping the model’s distribution rather than as a learning method or strategy. Consequently, the effect of distribution shaping is largely achieved within a very small number of training steps, leading to a decoupling between the continued decrease in EM training loss and improvements in mathematical reasoning performance.
@@ 3.5 Sampling Temperature in Training
Figure 3: The impact of generation temperature during EM training on the average performance of the trained model across six reasoning datasets. The results in the figure are obtained by repeating the experiments with the same training parameters using 16 different seeds to reduce randomness.
As shown in Figure 3, the average performance of the EM-trained model across six math reasoning benchmarks generally exhibits an upward trend as the generation temperature increases. The maximum of the average performance initially increases and then declines around a temperature of 0.5. Therefore, we prioritize the model trained at temperature 0.5 when reporting final results.
However, as shown in the figure, EM training exhibits significant randomness. The results in the figure are obtained by repeating experiments with 16 different random seeds under the same set of hyperparameters. It can be seen that, even with identical settings, the average scores across the six math reasoning benchmarks can differ by as much as a factor of two depending on the seed. Therefore, all the conclusions in this paper are based on at least 16 repeated experiments with different seeds. We also advocate that future research should focus on reducing the stochasticity of EM training.
@@ 3.6 Sampling Temperature in Evaluate Generation
Figure 4: The impact of generation temperature during evalutating on the average performance of the trained model across six reasoning datasets. The results in the figure are obtained by repeating the experiments with the same training parameters using 16 different seeds to reduce randomness.
As shown in Figure 4, we observe a striking pattern in the EM-trained model’s response to varying sampling temperatures during generation. Specifically, as the sampling temperature increases during evaluation, the model’s average performance across six math reasoning benchmarks consistently decreases. This trend is in sharp contrast to that of reinforcement learning (RL)-trained models shown in Figure 4, where higher sampling temperatures often improve performance.
@@
Together with our analysis in Section 3.3, we hypothesize that EM training systematically reshapes the model’s logit distributions to become increasingly right-skewed. This process reinforces confidence in already high-probability tokens, effectively concentrating the probability mass on semantically coherent and correct options. As a result, greedy decoding—which deterministically selects the most probable token—becomes a highly effective strategy post-EM fine-tuning.
In contrast, RL-based methods adjust token probabilities based on external reward signals tied to ground-truth outputs. This often promotes the relative ranking of previously low-probability (tail) tokens. Even after reranking, these tokens tend to occupy intermediate positions in the probability distribution, requiring higher temperatures during sampling to be selected. Consequently, RL-trained models exhibit the opposite trend: performance improves with higher sampling temperatures, as seen in Figure 4.
@@ 3.7 EM on Various Models
In Table 3, we report the peak performance of 1-shot EM after minimal training steps on various base models. Remarkably, with only a single exemplar and minimal optimization, EM consistently boosts inference accuracy across all models on tasks MATH500, Minerva Math, Olympiad Bench, and AMC23—without sacrificing general-purpose capabilities. For instance, on the Qwen2.5-Math-7B model, 1-shot EM yields impressive improvements of 25.8 points on MATH500 (from 53.0 to 78.8), 24.3 points on Minerva Math (from 11.0 to 35.3), 22.5 points on Olympiad Bench (from 17.2 to 39.7), and 26.2 points on AMC23 (from 44.1 to 70.3). This demonstrates EM’s ability to substantially enhance reasoning performance with virtually no extra data or compute. Consequently, EM emerges as a compelling lightweight alternative to RL across diverse model bases.
@@
3.8 EM Before/After RL
Figure 5: The blue curve on the left represents the average performance across six mathematical reasoning benchmarks of the model trained with RL during the EM phase as training steps progress, while the curve on the right shows the performance of the model trained with EM during the RL phase. The results in the figure are obtained by repeating the experiments with the same training parameters using 16 different seeds to reduce randomness.
Figure 5 shows a clear asymmetry: applying entropy minimization (EM) after reinforcement learning (RL) leads to a steady decline in performance across six math benchmarks, whereas applying RL after EM yields consistent gains. This suggests that EM exacerbates the distributional distortions introduced by RL, reinforcing the “alignment tax” of RLHF.
This aligns with prior work showing that RL is most effective when preceded by supervised fine-tuning (SFT), while applying SFT or entropy-based methods after RL often harms performance. In our case, EM after RL locks in narrow, overconfident output modes, while EM before RL enhances reasoning and allows RL to refine outputs without degrading diversity or accuracy.
@@ 3.9 Training Learning Rate
Figure 6: The impact of learning rate during EM training on the average performance of the trained model across six reasoning datasets. The results in the figure are obtained by repeating the experiments with the same training parameters using 16 different seeds to reduce randomness.
As shown in the Figure 6, a learning rate of 2e-5 demonstrates a significant advantage. We therefore recommend using 2e-5 as the default baseline setting for EM training.
3.10 1-shot vs. Multi-shot
Model MATH500 Minerva Math Olympiad Bench AMC23 AVG
LLaMA-3.1-8B 49.0 21.3 15.4 25.6 23.6
+ EM 49.2 23.2 13.8 25.9 24.0
+ EM 1-shot 42.6 21 17 37.8 24.3
Qwen2.5-7B 59.8 10.3 29.2 40.3 29.6
+ EM 61.8 18.4 31.1 43.4 33.5
+ EM 1-shot 67.4 22.1 33.6 45.6 37.3
Qwen2.5-7B-Instruct 77.0 37.1 37.9 51.9 43.12
+ EM 74.6 34.2 36.4 48.4 41.14
+ EM 1-shot 76.4 36.4 39.3 54.1 44.5
SimpleRL-Zoo 76.8 30.9 39.4 55.3 44.14
+ EM 77.2 29.0 35.7 55.9 41.8
+ EM 1-shot 73.0 24.6 35.3 45.9 39.1
Qwen2.5-Math-7B 53.0 11.0 17.2 44.1 27.7
+ EM 68.0 18.8 33.2 61.6 41.0
+ EM 1-shot 78.8 35.3 39.7 70.3 48.2
As shown in Table 1, EM training with just one or two examples achieves an average score on six math benchmarks that is on par with a dataset comprising thousands of examples, and even yields improvements, respectively. Data in Table 1 further demonstrate that EM using a minimal exemplar set delivers superior generalization and stronger performance across diverse downstream tasks.
Furthermore, as depicted in Figure 2, EM exhibits RL-like post-saturation generalization behavior: the 1-shot EM loss saturates at step 10, highlighting its remarkably fast convergence—just one example is sufficient to achieve significant gains; yet, model performance continues to improve steadily as training proceeds, implying that additional training steps beyond initial convergence can further enhance generalization once the model has fully assimilated the exemplar.
Notably, in the late stage of training, EM loss and model performance become decoupled—further loss reduction does not translate into performance gains and may even incur a slight degradation. We attribute this phenomenon to EM’s “over-confidence” effect, wherein the model amplifies its own prior biases, ultimately leading to a performance penalty.
4 Discussion and Future Work

### v2

@@ Abstract
We trained 13,440 large language models and found that entropy minimization requires only a single unlabeled data and 10 steps optimization to achieve performance improvements greater than those obtained using thousands of data and carefully designed rewards in rule-based reinforcement learning. This striking result may prompt a rethinking of post-training paradigms for large language models. Our code is avaliable at https://github.com/zitian-gao/one-shot-em.
@@ 1 Introduction
The post-training phase of large language models (LLMs) has advanced rapidly [22, 21, 9, 20, 3, 24], with models like DeepSeek-R1 [4], Kimi-K1.5 [19], and OpenAI o-series [13, 14] demonstrating remarkable reasoning abilities. However, preparing for Reinforcement Learning (RL) is never an easy task. It often requires a large amount of high-quality ground truth labeled data, along with the careful design of rule-based rewards to maximize advantage signals and prevent reward hacking. In contrast, Entropy Minimization (EM) is entirely unsupervised. We used the EM method to train 13,440 large language models in order to eliminate randomness in training as much as possible and ensure that the experimental results and observed patterns are reliable. Our rigorous study demonstrates that using just a single piece of unlabeled data, the performance already surpasses that of traditional RL. Moreover, they typically converge within just 10 training steps, which is significantly faster than the thousands of steps often required for RL. EM is based on two direct and simple assumptions: (1) The sampling process in generation of large language models is inherently stochastic, (2) Correct answers generally have lower entropy than incorrect ones. Our study reveals that EM and RL share the same goal: unlocking the pretrained model’s latent potential without adding new knowledge[11]. Both rely on a process we call “token reranking” to maximize the model’s performance. We find that entropy minimization has the capacity to rival RL in the post-training phase.
@@
  • We conduct an in-depth analysis of the effectiveness of One-shot EM, making it a highly reasonable approach. We find that it shares many core properties with RL, yet drives model behavior in the opposite direction when viewed through the lens of logits shift.
@@
  • We reveal that EM is a distribution shaping tool rather than a learning method by analyzing the inconsistency in loss-reasoning curve 3.4 and the logit shift 3.3 effect of EM.
@@ 2.2 Data Selection
Entropy minimization (EM) relies on the premise that the model’s predictive uncertainty can serve as a meaningful training signal. However, not all input prompts are equally informative in this regard. As noted in prior work [16, 20], certain prompts elicit deterministic behavior from the model (e.g., always correct or always incorrect), yielding limited gradient information under entropy-based objectives.
@@
This variance quantifies the inconsistency of the model’s predictions for a given input. A low variance indicates either high confidence in correctness (near-perfect success) or high confidence in failure (uniformly wrong), both of which are suboptimal for entropy minimization, as they lead to low-entropy posteriors that cannot be further improved.
@@
Intuitively, data with high pass@k variance suggests that the model’s response distribution is straddling the decision boundary—sometimes correct, sometimes not—indicating a broad or multimodal predictive distribution. These are precisely the regions where entropy minimization is most impactful: it encourages the model to concentrate its probability mass on a consistent and (ideally) correct reasoning trajectory.
@@ 3.1 Experimental Setting
Temperature Batch Size Learning Rate Max Response Len
0.5 64 $2\times 10^{-5}$ 2048
Table 1: Training hyper-parameters
We implemented the overall training process of EM based on Acclerate [7]. We selected 1 piece of data from the dataset as prompt. Since it is an unsupervised method, we do not need any data labels. We directly train the model for only 10 steps with a constant learning rate of $2\times 10^{-5}$, a temperature parameter of 0.5, and a batch size of 64.  [^1] The reason why only 10 steps are sufficient will be explained in detail in Section 3.4.
@@ 3.2 Main Result
During training, the model repeatedly reinforces its own confidence, causing the logits distribution to gradually shift to the right—a phenomenon that will also be discussed in detail in Section 3.3.
@@ 3.3 Logits Shift
We sample 20 prompts from the NuminaMath[10] dataset and generate responses using four different models (Qwen2.5-Math-7B, Qwen2.5-Math-7B-EM, Qwen2.5-Math-7B-RL, Qwen2.5-Math-7B-EM-RL). Each model generates 20 responses, resulting in a total of $4\times 20=80$ outputs. For each output, we extract the logits, defined as the unnormalized scores $z_{i}\in\mathbb{R}$ produced by the model before applying the softmax function.
@@
In contrast, models trained with RL show a pronounced reduction in logits skewness, indicating a leftward shift in the distribution. We hypothesize that this is due to the influence of ground-truth signals during RL training. Specifically, RL appears to perform re-ranking that suppresses tokens with high predicted probability but low alignment with ground truth, thereby reducing their rank and shifting the overall distribution leftward. Similarly, models trained with EM followed by RL exhibit the same pattern: after RL, the skewness of the model’s logits decreases from 1.54 (after EM) to 1.47, aligning with this pattern.
We refer to this phenomenon as logits shift. A rightward logits shift is beneficial to the generation and sampling process of large language models, as the model primarily samples from high-probability tokens. Shifting the logits to the right increases the number of candidate tokens and expands high-probability paths, thereby enhancing model capability. In contrast, a leftward logits shift is considered harmful to the generation process of large language models, as it reduces the number of high-probability paths during sampling—opposite to the effect of a rightward shift—thus diminishing the model’s overall performance. Therefore, the rightward logits shift induced by EM is preferable to the leftward shift caused by RL.
@@
Figure 2: The left y-axis represents the EM training loss, while the right y-axis shows the average score across four reasoning benchmarks. It can be observed that the training loss converges rapidly, whereas the average score peaks around step 10 and then begins to decline. The results in the figure are obtained by repeating the experiments with the same training hyper-parameters using 16 different seeds to reduce randomness.
As shown in Figure 2, the loss drops to a relatively low level around step 10, and the model’s performance on mathematical reasoning reaches its peak. However, unexpectedly, as the EM training loss continues to decrease beyond step 10, the mathematical reasoning performance begins to decline. In conjunction with the findings presented in Section 3.3, we argue that EM functions primarily as a tool for shaping the model’s distribution rather than as a learning method or strategy. Consequently, the effect of distribution shaping is largely achieved within a very small number of training steps, leading to a decoupling between the continued decrease in EM training loss and improvements in mathematical reasoning performance.
@@ 3.5 Sampling Temperature in Training
Figure 3: The impact of generation temperature during EM training on the average performance of the trained model across four reasoning datasets. The results in the figure are obtained by repeating the experiments with the same training hyper-parameters using 16 different seeds to reduce randomness.
As shown in Figure 3, the average performance of the EM-trained model across four math reasoning benchmarks generally exhibits an upward trend as the generation temperature increases. The maximum of the average performance initially increases and then declines around a temperature of 0.5. Higher temperatures lead to better average reasoning ability, while moderate temperatures (e.g., 0.5) result in greater performance variance, thereby creating opportunities for higher peak performance. Therefore, we prioritize the model trained at temperature 0.5 when reporting final results.
However, as shown in the figure, EM training exhibits significant randomness. The results in the figure are obtained by repeating experiments with 16 different random seeds under the same set of hyperparameters. It can be seen that, even with identical settings, the average scores across the four math reasoning benchmarks can differ by as much as a factor of two depending on the seed. Therefore, all the conclusions in this paper are based on at least 16 repeated experiments with different seeds. We also advocate that future research should focus on reducing the stochasticity of EM training.
@@ 3.6 Sampling Temperature in Evaluate Generation
Figure 4: The impact of generation temperature during evalutating on the average performance of the trained model across four reasoning datasets. The results in the figure are obtained by repeating the experiments with the same training hyper-parameters using 16 different seeds to reduce randomness.
As shown in Figure 4, we observe a striking pattern in the EM-trained model’s response to varying sampling temperatures during generation. Specifically, as the sampling temperature increases during evaluation, the model’s average performance across four math reasoning benchmarks consistently decreases. This trend is in sharp contrast to that of Reinforcement Learning (RL)-trained models shown in Figure 4, where higher sampling temperatures often improve performance.
@@
Together with our analysis in Section 3.3, we hypothesize that EM training systematically reshapes the model’s logits distributions to become increasingly right-skewed. This process reinforces confidence in already high-probability tokens, effectively concentrating the probability mass on semantically coherent and correct options. As a result, greedy decoding—which deterministically selects the most probable token—becomes a highly effective strategy post-EM fine-tuning.
In contrast, RL adjust token probabilities based on external ground-truth reward. This often promotes the relative ranking of previously low-probability (tail) tokens. Even after reranking, these tokens tend to occupy intermediate positions in the probability distribution, requiring higher temperatures during sampling to be selected. Consequently, RL-trained models exhibit the opposite trend: performance improves with higher sampling temperatures, as seen in Figure 4.
3.7 EM Before/After RL
Figure 5: The blue curve on the left represents the average performance across four mathematical reasoning benchmarks of the model trained with RL during the EM phase as training steps progress, while the curve on the right shows the performance of the model trained with EM during the RL phase. The results in the figure are obtained by repeating the experiments with the same training hyper-parameters using 16 different seeds to reduce randomness.
Figure 5 shows a clear asymmetry: applying Entropy Minimization (EM) after Reinforcement Learning (RL) leads to a steady decline in performance across four math benchmarks, whereas applying RL after EM yields consistent gains. This suggests that EM exacerbates the distributional distortions introduced by RL, reinforcing the “alignment tax” of RL.
This aligns with prior work showing that RL is most effective when preceded by Supervised Fine-Tuning (SFT) [2, 6], while applying SFT or entropy-based methods after RL often harms performance. In our case, EM after RL locks in narrow, overconfident output modes, while EM before RL enhances reasoning and allows RL to refine outputs without degrading diversity or accuracy.
@@ 3.8 EM on Various Models
In Table 3, we report the peak performance of 1-shot EM after minimal training steps on various base models. Remarkably, with only a single exemplar and minimal optimization, EM consistently boosts inference accuracy across all models on tasks MATH500, Minerva Math, Olympiad Bench, and AMC23. For instance, on the Qwen2.5-Math-7B model, 1-shot EM yields impressive improvements of 25.8 points on MATH500 (from 53.0 to 78.8), 24.3 points on Minerva Math (from 11.0 to 35.3), 22.5 points on Olympiad Bench (from 17.2 to 39.7), and 26.2 points on AMC23 (from 44.1 to 70.3). This demonstrates EM’s ability to substantially enhance reasoning performance with virtually no extra data or compute. Consequently, EM emerges as a compelling lightweight alternative to RL across diverse model bases.
@@
3.9 1-shot vs. Multi-shot
Model MATH500 Minerva Math Olympiad Bench AMC23 Avg.
LLaMA-3.1-8B 49.0 21.3 15.4 25.6 27.8
+ EM multi-shot 49.2 23.2 13.8 25.9 28.0
+ EM 1-shot 42.6 21 17 37.8 29.6
Qwen2.5-7B 59.8 10.3 29.2 40.3 34.9
+ EM multi-shot 61.8 18.4 31.1 43.4 38.7
+ EM 1-shot 67.4 22.1 33.6 45.6 42.2
Qwen2.5-7B-Instruct 77.0 37.1 37.9 51.9 50.9
+ EM multi-shot 74.6 34.2 36.4 48.4 48.4
+ EM 1-shot 76.4 36.4 39.3 54.1 51.5
SimpleRL-Zoo 76.8 30.9 39.4 55.3 50.6
+ EM multi-shot 77.2 29.0 35.7 55.9 49.5
+ EM 1-shot 73.0 24.6 35.3 45.9 44.7
Qwen2.5-Math-7B 53.0 11.0 17.2 44.1 31.3
+ EM multi-shot 68.0 18.8 33.2 61.6 45.4
+ EM 1-shot 78.8 35.3 39.7 70.3 56.0
As shown in Table 2, EM training with just one or two examples achieves an average score on four math benchmarks that is on par with a dataset comprising thousands of examples, and even yields improvements, respectively. Data in Table 2 further demonstrate that EM using a minimal exemplar set delivers superior generalization and stronger performance across diverse downstream tasks.
Furthermore, as depicted in Figure 3.4, EM exhibits RL-like post-saturation generalization behavior: the 1-shot EM loss saturates at step 10, highlighting its remarkably fast convergence—just one example is sufficient to achieve significant gains.
Notably, in the late stage of training, EM loss and model performance become decoupled—further loss reduction does not translate into performance gains and may even incur a slight degradation. We attribute this phenomenon to EM’s "over-confidence" effect, wherein the model amplifies its own prior biases, ultimately leading to a performance penalty.
@@ 3.10 Training Learning Rate
Figure 6: The impact of learning rate during EM training on the average performance of the model at step 10 across four reasoning datasets. The results in the figure are obtained by repeating the experiments with the same training hyper-parameters using 16 different seeds to reduce randomness.
As shown in Figure 6, a learning rate of $2\times 10^{-5}$ demonstrates a significant advantage. Since EM typically converges and reaches peak performance within just 10 steps, an excessively large learning rate may cause severe and rapid overconfidence, while a smaller learning rate can slow down convergence. To improve training efficiency, we therefore recommend using $2\times 10^{-5}$ as the default baseline setting for EM training.
4 Future Work